BIBTEX(1) BIBTEX(1) NAME bibtex−makeabibliographyfor(La)TeX SYNOPSIS bibtex[-min-crossrefs=number][-terse][auxname] DESCRIPTION Thismanualpageisnotmeanttobeexhaustive.Thecompletedocumentationforthisversionof TEXcanbefoundintheinfofileormanualWeb2C:ATeXimplementation. BIBTEXreadsthetop-levelauxiliary(.aux)filethatwasoutputduringtherunningoflatex(1)or tex(1)andcreatesabibliography(.bbl)filethatwillbeincorporatedintothedocumentonsubsequentrunsofLATEXorTEX.Theauxnameonthecommandlinemustbegivenwithoutthe.aux extension.Ifyoudon’tgivetheauxname,theprogrampromptsyouforit. BIBTEXlooksup,inbibliographicdatabase(.bib)filesspecifiedbythe\bibliographycommand, theentriesspecifiedbythe\citeand\nocitecommandsintheLATEXorTEXsourcefile.Itformatstheinformationfromthoseentriesaccordingtoinstructionsinabibliographystyle(.bst)file (specifiedbythe\bibliographystylecommand,anditoutputstheresultstothe.bblfile. TheLATEXmanualexplainswhataLATEXsourcefilemustcontaintoworkwithBIBTEX. AppendixBofthemanualdescribestheformatofthe.bibfiles.The‘BIBTEXing’document describesextensionsanddetailsofthisformat,anditgivesotherusefulhintsforusingBIBTEX. OPTIONS The-min-crossrefsoptiondefinestheminimumnumberofcrossrefrequiredforautomatic inclusionofthecrossref’dentryonthecitationlist;thedefaultistwo.Withthe-terseoption, BIBTEXoperatessilently.Withoutit,abannerandprogressreportsareprintedonstdout. ENVIRONMENT BIBTEXsearchesthedirectoriesinthepathdefinedbytheBSTINPUTSenvironmentvariablefor .bstfiles.IfBSTINPUTSisnotset,itusesthesystemdefault.For.bibfiles,itusestheBIBIN-PUTSenvironmentvariableifthatisset,otherwisethedefault.Seetex(1)forthedetailsofthe searching. IftheenvironmentvariableTEXMFOUTPUTisset,BIBTEXattemptstoputitsoutputfilesinit, iftheycannotbeputinthecurrentdirectory.Again,seetex(1).Nospecialsearchingisdonefor the.auxfile. FILES *.bstBibliographystylefiles. btxdoc.tex ‘‘BIBTEXing’’−LATEXabledocumentationforgeneralBIBTEXusers btxhak.tex ‘‘DesigningBIBTEXStyles’’−LATEXabledocumentationforstyledesigners btxdoc.bib databasefileforthosetwodocuments xampl.bib databasefilegivingexamplesofallstandardentrytypes btxbst.doc templatefileanddocumentationforthestandardstyles Allthosefilesshouldbeavailablesomewhereonyoursystem. Thehostmath.utah.eduhasavastcollectionof.bibfilesavailableforanonymousftp,including referencesforallthestandardTEXbooksandacompletebibliographyforTUGboat. Web2C7.32x 24November1994 1 BIBTEX(1) BIBTEX(1) SEEALSO latex(1),tex(1). LeslieLamport,LATEX−ADocumentPreparationSystem,Addison-Wesley,1985,ISBN 0-201-15790-X. AUTHOR OrenPatashnik,StanfordUniversity.Thismanpagedescribestheweb2cversionofBIBTEX. OtherportsofBIBTEX,suchasDonaldKnuth’sversionusingtheSunPascalcompiler,donot havethesamepathsearchingimplementation,orthecommand-lineoptions. Web2C7.32x 24November1994 2